Find the magnitude and arguments of: a) Z1 7 - 12i b) Z2=-1 i

please give correct answer 100% please make sure hand writing is clear and easy to read. please help me explain clearly the answer given because i need to understand.
Answer #1

a) 2,=7-12) a=4, b=-12 J use o Magnitude = r 737 6125 = V 49 +144 = 13.892 Argument is aso, bo use o 0,= tant (2) = -59.743 b